The sum of four numbers is 771. The ratio of the first number to the second number is 2:3. The ratio of the second number to the third number is 5:4. The ratio of the third number to the fourth number is 5:6. What is the second number?
Answer by edjones(8007)

Let the 1st number = x
x + 3x/2 + 3x/2 * 4/5 + 6x/5 * 6/5
x + 3x/2 + 6x/5 + 36x/25 = 771
50x+75x+60x+72x=38550 Multiply each side by 50.
257x=38550
x=150
3x/2=225 the 2nd number
.
Ed